OpenStudy (anonymous):

Simplify (4^3√2+2√3)-(2^3√16-4√3)

OpenStudy (whpalmer4):

\[(4^3\sqrt{2}+2\sqrt{3}) - (2^3\sqrt{16}-4\sqrt{3})\]?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Well, I would first evaluate the 4^3 and 2^3

OpenStudy (anonymous):

then?

OpenStudy (whpalmer4):

What do you get?

OpenStudy (whpalmer4):

I know, I said "I would" but I meant "you should" :-)

OpenStudy (whpalmer4):

Next, I would look at the square roots and see if any of them can be directly evaluated, or factored. One of them can be.

OpenStudy (anonymous):

6√3 right?

OpenStudy (whpalmer4):

Um, no. What did you get for 4^3 and 2^3?

OpenStudy (anonymous):

(4^3√2+2√3)-(2x2^3√2-4√3)

OpenStudy (whpalmer4):

No, what is the value of 4^3 and 2^3? What is 4*4*4? What is 2*2*2?